### v1

# EchoSight: Advancing Visual-Language Models with Wiki Knowledge

Abstract
Knowledge-based Visual Question Answering (KVQA) tasks require answering questions about images using extensive background knowledge. Despite significant advancements, generative models often struggle with these tasks due to the limited integration of external knowledge. In this paper, we introduce EchoSight, a novel multimodal Retrieval-Augmented Generation (RAG) framework that enables large language models (LLMs) to answer visual questions requiring fine-grained encyclopedic knowledge. To strive for high-performing retrieval, EchoSight first searches wiki articles by using visual-only information, subsequently, these candidate articles are further reranked according to their relevance to the combined text-image query. This approach significantly improves the integration of multimodal knowledge, leading to enhanced retrieval outcomes and more accurate VQA responses. Our experimental results on the Encyclopedic VQA and InfoSeek datasets demonstrate that EchoSight establishes new state-of-the-art results in knowledge-based VQA, achieving an accuracy of 41.8% on Encyclopedic VQA and 31.3% on InfoSeek.
EchoSight: Advancing Visual-Language Models with Wiki Knowledge
Yibin Yan1,2           Weidi Xie1,2 1Shanghai Jiao Tong University        2Shanghai AI Lab https://go2heart.github.io/echosight
Figure 1: For visual questions such as “When was the 1st ascent of this mountain?”, visual-only search methods consider image similarity only, ignoring the textual details of the accompanying article. By incorporating multimodal reranking, the correct entry, accounting for both visual and textual information, can be accurately identified.
1 Introduction
Visual Question Answering (VQA) addresses the challenge of enabling machines to understand and respond to questions about visual content, typically images or videos. Broadly, this task can be divided into two categories: standard VQA Antol et al. (2015); Goyal et al. (2017) with questions that can be answered directly from the visual content, for example, counting objects, identifying colors, or recognizing simple actions, which rely solely on commonsense and information present in the image; and knowledge-based VQA Marino et al. (2019); Schwenk et al. (2022); Mensink et al. (2023); Chen et al. (2023) requiring additional context or external knowledge, such as historical facts, detailed object properties, or specific situational contexts not evident in the visual content.
Addressing these two types of questions presents different challenges for VQA systems. Questions that draw answers directly from visual content demand robust image understanding capabilities, encompassing tasks such as object detection, scene recognition, and spatial reasoning. Conversely, questions requiring external knowledge call for additional mechanisms to access and integrate information from external sources. In this paper, we focus on the latter type of visual question answering, by building a retrieval-augmented multimodal system, that enables searching an external knowledge base for more nuanced understanding and accurate responses.
Despite the recent accomplishments in developing Visual-language Models (VLMs) Achiam et al. (2023); Gemini Team et al. (2023); Abdin et al. (2024); Liu et al. (2024), knowledge-based VQA remains challenging. This complexity primarily stems from two aspects. (i) Existing VLMs struggle to adequately encode all essential knowledge, due to its limited model capacity, and infrequent inclusion of encyclopedic, long-tail information in their training data Mensink et al. (2023). (ii) The visual component of the questions often provides limited help in addressing the queries, as establishing a meaningful connection between entity knowledge and visual attributes can be difficult. For example, an image of a church alone does not reveal information about its construction date.
In this paper, we introduce EchoSight, a novel retrieval-augmented vision-language system designed for knowledge-based visual question answering. EchoSight employs a dual-stage search mechanism that integrates a retrieval-and-reranking process with the Retrieval Augmented Generation (RAG) paradigm. Initially, the system performs a visual-only retrieval from an external knowledge base, to effectively narrow the knowledge search space, only focusing on candidates that are closely align with the visual context of the reference image. In the subsequent multimodal reranking stage, the system refines the candidates ranking by incorporating both the reference image and the textual query. This approach guarantees that the selected results are pertinent not only visually, but also contextually to the multimodal query. After acquiring the most relevant information through this coarse-to-fine grained search, our model generates the precise answer to the posed question.
Overall, we present three contributions: First, we propose a multimodal retrieval-augmented generation framework, termed as EchoSight, that enables LLMs to answer visual questions that require fine-grained encyclopedic knowledge; Second, we adopt a retrieval-and-reranking scheme to improve retrieval performance, specifically, it first searches images with visual-only information, and then conduct a fine-grained multimodal reranking on the candidates; Third, we conduct thorough experiments on both Encyclopedic VQA Mensink et al. (2023) and InfoSeek Chen et al. (2023) benchmarks, EchoSight demonstrates state-of-the-art performance on both benchmarks, significantly outperforming existing VLMs or other retrieval-augmented architectures.
Figure 2: The overall view of our proposed EchoSight. (i) Given a visual question with an image, the retriever searches the reference image in the knowledge base for top $k$ similar images to get their corresponding Wikipedia Entries. (ii) After changing the granularity to sections, all the sections of retrieved entries are then reranked with the maximum pairwise similarity of their textual embeddings and the reference image+question’s Q-Former query tokens. (iii) The top reranked section will be utilized as RAG prompt for the LLM to generate the ultimate answer.
2 Method
This section starts with the problem formulation of retrieval-augmented VQA (Sec. 2.1), followed by detailing the retrieval-and-reranking module in EchoSight (Sec. 2.2), and finally the answer generation module (Sec. 2.3).
2.1 Problem Formulation
Given a reference image, and question of free-form texts, our goal is to construct a visual question answering system, that can benefit from the access of an external knowledge base. In our case, this is a million-scale dataset of entity articles and their corresponding images from Wikipedia webpage, i.e., =$\mathcal{B}=\{(a_{1},I_{1}),\dots,(a_{n},I_{n})\}$.
The overall architecture of our proposed method, EchoSight, is illustrated in Figure 2. It consists of four main components: an external knowledge base (KB), a retriever, a reranker, and an answer generator. (i) The process begins with the retriever, which utilizes the reference image to filter and extract relevant KB entries with similar images; (ii) Next, the reranker takes these candidate entries and employs their textual contents to have them reranked, based on their relevance to both the reference image and the textual question; (iii) Finally, the reranked KB entries are fed into the answer generator to produce the final answer.
2.2 Retrieval and Reranking
The goal of this stage is to identify relevant entries from a large-scale external knowledge base using the given reference image and question. We employ a two-stage procedure: first, a visual-only search identifies candidates that are visually similar to the query image. Subsequently, a multimodal reranking process evaluates both visual and textual information to reorder the retrieved entries. This ensures that the most pertinent article entry can be ranked at the top, facilitating efficient and accurate answer generation.
Visual-only Search. Given the extensive size of the knowledge base, potentially encompassing millions of image-article pairs, optimizing the efficiency of the image search process is critical. To achieve this, we transform all images into vectors and utilize the cosine similarity metric to assess their proximity to a reference image.
$S_{\Omega}=\Biggl{\{}s_{i}=\biggl{\langle}\frac{v_{r}}{||v_{r}||}\cdot\frac{v_{i}}{||v_{i}||}\biggr{\rangle},i=1,\dots,n\Biggl{\}},$
where $v_{r}=\Phi_{\text{vis}}(I_{\text{ref}})$ and $v_{i}=\Phi_{\text{vis}}(I_{i})$ denote the visual embedding for reference image and database image, respectively, computed by a pre-trained visual encoder. We employ the FAISS libraryDouze et al. (2024) for vector search, and keep the top $k$ best-matched images and their corresponding wiki article entries from the knowledge base, i.e., $\mathcal{E}_{v}=\{(a_{1},I_{1}),\dots,(a_{k},I_{k})\},k\ll n$.
Multimodal Reranking. After initially filtering the candidates based on visual similarities, the reranker module integrates both textual and visual inputs from the multimodal query and the top $k$ retrieved Wikipedia article entries. This stage aims to prioritize entries that are most pertinent to the question, ensuring the articles with highest relevancy are ranked at the top.
Specifically, we employ the Q-Former Li et al. (2023b) architecture to extract multimodal information from the reference image and textual question, resulting $32$ query tokens.
$z_{m}^{i}=\text{Q-Former}\left(I_{\text{ref}},Q\right)^{i},$
where $z_{m}^{i}$ denotes the $i$th query token embedding of the reference image $I_{ref}$ and textual question $Q$.
On the candidates side, we break each of the wiki articles into sections, with each section prefixed by the article’s title, for example, $a_{i}=\{\text{sec}_{1}^{i},\text{sec}_{2}^{i},\dots,\text{sec}_{p}^{i}\}$, and further encode them with Q-Former’s text encoder. We initialize the Q-Former with BLIP-2’s weights and fine-tunes with all parameters open except the visual encoder.
The reranking score for each section is calculated as follows:
$S_{r}^{\text{sec}}=\max_{1\leq i\leq N_{q}}\left(\text{sim}(z_{m}^{i},z_{s}^{\text{sec}})\right),$
where $S_{r}^{\text{sec}}$ is the reranking score for section “sec”, determined using the Q-Former’s Image-to-Text Correspondence (ITC) method. This method computes the highest pairwise similarity between each multimodal query token embedding $z_{m}^{i}$ from the reference image and question pair, and the $[CLS]$ token embedding of a Wikipedia article section $z_{s}^{\text{sec}}$. $N_{q}$ denotes the number of query tokens.
In the final step of multimodal reranking, the reranker combines the visual similarity score from the previous stage and the reranking score into a weighted sum:
$\text{sec}_{vl}=\operatorname*{arg\,max}_{\text{sec}\in\mathcal{E}_{v}}\left(\alpha\cdot S_{v}^{\text{sec}}+(1-\alpha)\cdot S_{r}^{\text{sec}}\right),$
where $\text{sec}_{vl}$ refers to the highest-ranked entry section produced by the reranker, $\alpha$ is a weight parameter that balances the visual similarity score $S_{v}^{\text{sec}}$ and the reranking score $S_{r}^{\text{sec}}$. Note that $S_{v}^{sec}$ is calculated in the visual-only search stage using the best-matched image from the wiki entry to which $sec$ belongs.
Reranker Training. Here, we implement hard negative sampling within a contrastive learning framework. Specifically, hard negative samples are specifically selected from examples that are visually similar yet contextually distinct, i.e., the initial visual-only retrieval efforts were unsuccessful. With such training, the reranker is thus forced to select the most relevant articles for the multimodal queries, enhancing the overall accuracy and effectiveness of the system Robinson et al. (2021).
The training objective of the reranker is given as follows:
$\mathcal{L}=-\log\frac{\exp(\max_{1\leq i\leq N_{q}}\text{sim}(z_{m}^{i},z_{s}))}{\sum_{j=1}^{N}\exp(\max_{1\leq i\leq N_{q}}\text{sim}(z_{m}^{i},z_{s}^{j}))},$
where $z_{s}$ denotes the positive section embedding, and $N$ is the total number of samples including both positive and negatives.
2.3 Answer Generation with LLMs
Once the relevant entries are identified from the knowledge base, large language models (LLMs) will integrate such information to answer the questions, i.e., $A=\text{LLM}(\text{sec}_{vl},Q)$, where the off-the-shelf LLM acts as an answer generator, $\text{sec}_{vl}$ denotes the retrieved wiki article section, and $Q$ refers to the target question. Comparing to existing generative VLMs, such retrieval-augmented generation (RAG) Lewis et al. (2020), enables the model with the essential contextual knowledge, improving the system’s ability to handle complex questions that demand precise and detailed knowledge.
3 Experiments
3.1 Datasets
Encyclopedic VQA Mensink et al. (2023) contains 221k unique question and answer pairs each matched with (up to) 5 images, resulting in a total of 1M VQA samples. These images are derived from iNaturalist 2021 (iNat21) Van Horn et al. (2021) and Google Landmarks Dataset V2 (GLDv2) Weyand et al. (2020). The visual questions focus on the fine-grained categories and instances. There are single-hop and two-hop questions that require different reasoning steps in the dataset. Notably, the dataset provides a controlled knowledge base with 2M Wikipedia articles with images, ensuring all the questions can be answered if correct Wikipedia article is given. For our experiments on E-VQA, we consider the single-hop questions using the provided 2M knowledge base.
InfoSeek Chen et al. (2023) comprises 1.3M visual information-seeking questions, covering more than 11K visual entities from OVEN Hu et al. (2023a). InfoSeek provides a knowledge base with 100K Wikipedia articles with images. The questions of the dataset are diverse and the answers can be referenced from Wikipedia. There are a human-labeled 8.9K collection and an automated generated 1.3M collection in InfoSeek. Due to the unavailability of groundtruth for test split, we report evaluation results on the validation split. We note that, the original authors did not publicly release their knowledge base, we therefore filter a 100K knowledge base from E-VQA instead. We will release ours to the community for reproduction and future comparison.
3.2 Metrics
To evaluate the performance of our proposed retrieval-augmented QA model, we focus on two aspects, namely, retrieval and question answering. The retrieval results gauge the system’s capability to accurately retrieve relevant articles from a large-scale multimodal knowledge base, while the question answering results assess its holistic effectiveness in providing precise and correct answers to visual questions
Metrics for Retrieval. We utilize the standard metric Recall@K. Recall@K assesses whether the correct article entries appear among the top $k$ retrieved results. An article is considered correct only if its URL exactly matches the target URL, making our retrieval evaluation more stringent and precise compared to methods that only match the content of answers to the retrieved articles.
Metrics for Question Answering. Here, we follow the conventional practise, use different metrics depending on the considered datasets. For E-VQA dataset Mensink et al. (2023), we use the BEM (Balanced Evaluation Metric) score Zhang et al. (2019), while for the InfoSeek dataset Chen et al. (2023), we employ the VQA accuracy Goyal et al. (2017); Marino et al. (2019) and Relaxed Accuracy Methani et al. (2020); Masry et al. (2022). These metrics are chosen to align with the evaluation settings specific to each dataset.
Method Recall@K
K=1 K=5 K=10 K=20
Google Lens 47.4 62.5 64.7 65.2
CLIP I-T 3.3 7.7 12.1 16.5
EchoSight
w/o. Reranking 13.3 31.3 41.0 48.8
w. Reranking 36.5 47.9 48.8 48.8
Table 1: E-VQA retrieval experiments. While Google Lens can be recognized as a upperbound in E-VQA, CLIP I-T indicates the retrieval from the reference image to Wikipedia entry texts with CLIP Radford et al. (2021).
3.3 Implementation Details
The Retriever. We compute the visual embedding for the reference images and images from database with a frozen Eva-CLIP vision encoder (Eva-CLIP-8B) Sun et al. (2024). The pooled last-layer embedding are used as the features for computing cosine similarity between images, with FAISS library.
The Reranker. The reranking module is initialized with pre-trained BLIP-2 Li et al. (2023b) weights using the LAVIS Library Li et al. (2023a). The number of query tokens $N_{q}$ is 32 and weighting parameter $\alpha$ is $0.5$. Instead of using in-batch contrastive learning, we employ hard negative sampling, where each positive sample is paired with $N=24$ negative samples.
In practise, a positive sample is constructed using the evidence section text from the corresponding Wikipedia article. While for negative samples, we perform a visual-only search on the reference images. Knowledge base entries with images that fail to match the reference images ranked within the top $k$ are selected as negative samples. During training, we randomly sample sections from these negative entries as well as from the non-evidence sections of the positive entries. Note that, as only E-VQA dataset provides labeled evidence sections for all its training data, we train the reranker on this dataset, and directly use it on InfoSeek in a zero-shot manner.
We adopt OneCycleLR Smith and Topin (2019) scheduler, with AdamW Loshchilov and Hutter (2018) optimizer. We use learning rate $10^{-4}$, batch size $6$, and the negative samples per example being $24$. For training the reranker module with 900K examples, 150K steps require 40 hours on 1 Nvidia A100 (80G).
The Answer Generator. We use Mistral-7B-Instruct-v0.2 Jiang et al. (2023) as the question generator for E-VQA and LLaMA-8B-Instruct AI@Meta (2024) for InfoSeek.
Method Recall@K
K=1 K=5 K=10 K=20
[…]
CLIP I-T 32.0 54.0 61.6 68.2
EchoSight
w/o. Reranking 45.6 67.1 73.0 77.9
w. Reranking 53.2 74.0 77.4 77.9
[…]
      Method       LLM       Retrieval       E-VQA       InfoSeek
      Google Lens       PaLM       KB Article       48.0       -
      Google Lens       PaLM       KB Section       48.8       -
      Vanilla       PaLM       -       19.7       1.0
      Mistral-7B       -       21.0       0.4
      LLaMA3-8B       -       18.7       2.4
      BLIP-2       Flan-T5XXL       -       12.6       12.5
      LLaVA-1.5       Vicuna-7B       -       16.3       9.5
      Wiki-LLaVA       Vicuna-7B       KB Section       21.8       28.9
[…]
      EchoSight
        w/o. Reranking       Mistral-7B | LLaMA3-8B [^1] The E-VQA accuracy is tested with Mistral-7B and InfoSeek accuracy is tested with LLaMA3-8B.       KB Article       19.4       27.7
        w. Reranking       Mistral-7B | LLaMA3-8B [^1] The E-VQA accuracy is tested with Mistral-7B and InfoSeek accuracy is tested with LLaMA3-8B.       KB Section       41.8       31.3
[…]
3.4 Results
In this section, we present experimental results on the E-VQA and InfoSeek benchmarks.
On Retrieval. The experiment results for the retrieval tasks across different configurations are detailed in Table 1 and Table 2. The CLIP I-T setting involves using CLIP for cross-modal similarity search, from the reference image to the Wikipedia article. The articles are represented as CLIP embedding of their title and descriptions. The ‘Google Lens’ refers to the approach used in Encyclopedic VQA Mensink et al. (2023), where Google Lens indexes billions of images from the Internet, not limited to Wikipedia, to find and return the most closely matching images along with an entity prediction. The best corresponding knowledge base entry identified by Google Lens is considered the result of its retrieval effort. Given its vast image index and capability to associate images with relevant entities, Google’s retrieval can be viewed as a upperbound in E-VQA retrieval.
From both tables, we can draw the observation that, our proposed reranking module has shown to significantly improve the retrieval performance, for example, it improves Recall@1 from 13.3% to 36.5% on E-VQA benchmark, 45.6% to 53.2% on InfoSeek, largely bridging the gap towards the ‘Google Lens’ upperbound.
[…]
It is clear that our proposed EchoSight (w. reranking) has outperform the prior works by a significant margin, even approaching the upperbound results reported by original E-VQA Mensink et al. (2023) benchmark, where two giant models are adopted, i.e., ‘Google Lens’ for knowledge retrieval, and PaLM as answer generation.
3.5 Ablation Study
For all experiments in ablation study, we use the E-VQA dataset. On the retrieval side, we conduct the following ablation studies: (i) to compare different visual backbones in retrieval module, (ii) to study the impact of reranking scope and (iii) to investigate the importance of hard negative sampling. On final answer generation, we carry out ablation studies on: (i) the impact of different language models, (ii) to experiment the answer generator under oracle retrieval results.
Backbone Recall@K
K=1 K=5 K=10 K=20
OpenAI-CLIP
w/o. Reranking 10.1 19.5 25.8 32.2
w. Reranking 23.8 31.4 32.1 32.2
Eva-CLIP
w/o. Reranking 13.3 31.3 41.0 48.8
w. Reranking 36.5 47.9 48.8 48.8
Table 4: Retrieval performance analysis on different vision backbones.OpenAI-CLIP is CLIP-ViT-Large Radford et al. (2021) and Eva-CLIP is Eva-CLIP-8B Sun et al. (2024) from BAAI. We both take the visual encoder’s last layer output as the image feature.
Impact of vision backbones. We assess the effect of different visual backbones on the retrieval stage, as detailed in Table 4. We compare the Vision Transformer (ViT) from EvaCLIP-8B Sun et al. (2024) with OpenAI’s CLIP-ViT-Large Radford et al. (2021). The EvaCLIP-8B’s ViT achieves a recall@20 of 48.8%, outperforming the CLIP-ViT-Large, which scored 32.2%. This substantial improvement is likely due to EvaCLIP-8B’s larger parameter size and more extensive training dataset, allowing it to develop more robust representations.
While the initial Recall@1 shows a modest difference between the two models (10% for CLIP-ViT-Large and 13% for EvaCLIP-8B), adopting our multimodal reranking significantly boosts performance, increasing Recall@1 to 23.8% and 36.5% for CLIP-ViT-Large and EvaCLIP-8B, respectively. This results in a marked 13% difference, underscoring the effectiveness of our approach, especially when combined with a more capable backbone.
Impact of reranking scope. The reranking scope refers to the number of candidates considered by the reranker module. Involving a higher reranking scope means calculating more embeddings during the reranking process. The reranking scope, which can be any number up to $k$, i.e., the total number of candidates returned by the retriever. As shown in Table 5, our reranker can consistently improve the results with increasing scope from Top-5 to Top-500, though it will significantly increase the computation cost, resulting in diminishing returns. Considering the balance of efficiency and quality, the scope of 20 candidate entries is used when reporting our final VQA accuracy on E-VQA and InfoSeek.
Scope Recall@K
K=1 K=5 K=10 K=20
Top 5 29.4 32.2 - -
Top 10 34.3 40.7 40.9 -
Top 20 36.5 47.9 48.8 48.8
Top 50 38.3 53.6 56.9 57.9
Top 100 38.8 55.9 60.8 63.0
Top 500 39.8 58.5 65.3 70.3
Table 5: The ablation study on impact of the reranking scope. Our reranker can consistently improve the results with increasing scope from Top-5 to Top-500.
Impact of hard negative sampling. The training strategy of the reranker module is critical for its performance. Rather than using randomly selected, irrelevant article entries, we employ a hard negative sampling during training, i.e., top negative candidates returned by the retriever. This approach ensures the reranker to be trained on more demanding examples, thereby improving its performance and robustness. The effects of different training strategies on reranking performance are detailed in Table 6.
Sampling Recall@K
K=1 K=5 K=10 K=20
EchoSight
w/o. Hard Neg 31.4 46.0 48.5 48.8
w. Hard Neg 36.5 47.9 48.8 48.8
Table 6: The ablation study of how sampling methods affect the overall retrieval-and-reranking performance.
Consistency of EchoSight across LLMs. The choice of LLMs influences the RAG paradigm greatly Shao et al. (2023); Hu et al. (2022). We compare PaLM Chowdhery et al. (2023), GPT-4 Achiam et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023) and LLaMA3-8B-Instruct AI@Meta (2024) as answer generators. Specifically, we provide them with same reranking results (KB entries). As shown in Table 7, the accuracy results are calculated with BEM Zhang et al. (2019) following Mensink et al. (2023). The results indicate that though better language models yield better scores, the overall performance across all tested language models is quite stable. This validates our method adapts well across modern language models.
LLMs GPT-4 PaLM Mistral LLaMA3
Accuracy 44.4 39.0 41.8 38.9
Table 7: The ablation study of impact of language models. The results are generated with the retrieval results of EchoSight with reranking scope 20.
Effect of oracle retrieval. Oracle retrieval indicates that the correct Wikipedia entry is always provided for generating the answer. As shown in Table 8, LLMs can almost flawlessly answer the question if oracle retrieval is provided.
     LLM      Retrieval      Accuracy
     PaLM      KB Title      31.0
     Mistral-7B      KB Title      29.4
     LLaMA3-8B      KB Title      32.0
     PaLM      KB Article      78.4
     Mistral-7B      KB Article      84.8
     LLaMA3-8B      KB Article      84.6
Table 8: The ablation study with VQA results on the effect of oracle retrieval.
4 Related Work
4.1 Visual Question Answering
Visual Question Answering (VQA) is the task of answering open-ended questions based on an image with natural language response. VQA tasks can be divided into two types: standard VQA and knowledge-based VQA.
Standard VQA. Datasets such as VQAv1 Antol et al. (2015), VQAv2 Goyal et al. (2017), and VizWiz Gurari et al. (2018) focus on questions that can be answered by analyzing the image content alone, without external information. These datasets typically cover questions about objects in the image, their attributes and other perceptual details that can be inferred from the visual input.
Knowledge-based VQA. The task involves questions that require information not present in the image. Pioneering datasets like OK-VQA Marino et al. (2019) and A-OKVQA Schwenk et al. (2022), which include questions needing knowledge beyond what is visually depicted, necessitate the integration of external world knowledge and commonsense reasoning. However, both datasets focus primarily on commonsense and general world knowledge, often neglecting more specialized or encyclopedic facts, and they do not provide external knowledge bases.
To fill this gap, datasets such as Encyclopedic VQA (E-VQA) Mensink et al. (2023) and InfoSeek Chen et al. (2023) have been developed. These datasets utilize Wikipedia as a knowledge base to provide detailed and specific information on various topics. E-VQA covers a wide range of topics like animals, plants, and landmarks, while InfoSeek focuses on info-seeking questions about various visual entities. These datasets require models to recognize visual entities and accurately retrieve and use relevant information from external sources Lerner et al. (2024); Caffagni et al. (2024); Lin et al. (2024).
4.2 Vison Language Models for VQA
Advances in Vision Language Models (VLMs) such as GPT-4V Achiam et al. (2023), Gemini Gemini Team et al. (2023), LLaVA Liu et al. (2024), and Phi-3-Vision Abdin et al. (2024) have demonstrated impressive capabilities in standard Visual Question Answering (VQA) tasks, exhibiting strong image analysis and accurate response generation Li et al. (2023d). However, these models encounter difficulties with knowledge-based VQA due to issues such as hallucination, where responses are generated based on nonexistent content and internal biases Li et al. (2023c), and the lack of efficient knowledge retrieval mechanisms which hampers the integration of external knowledge bases for reasoning Caffagni et al. (2024).
Recently, research has shifted towards retrieval-augmented generative systems. While Retrieval-Augmented Generation (RAG) has been well-established in Large Language Models (LLMs), its application in VLMs remains underexplored. Systems like KAT Gui et al. (2021), REVIVE Lin et al. (2022), and REVEAL Hu et al. (2023b) show promise for questions involving commonsense reasoning, yet they struggle with complex, knowledge-intensive tasks like Encyclopedic VQA (E-VQA) and Infoseek. These limitations stem from their restricted ability to fetch and incorporate precise information from extensive encyclopedic knowledge bases Mensink et al. (2023).
EchoSight addresses these issues through a novel two-stage process combining visual-only retrieval and multimodal reranking. This approach significantly enhances the alignment between retrieved textual knowledge and visual content, leading to improved performance on benchmarks such as Encyclopedic VQA and InfoSeek.
5 Conclusion
In this paper, we introduced EchoSight, a novel retrieval-augmented vision language system designed to address the challenges of knowledge-based Visual Question Answering (VQA). Our approach enhances the retrieval capabilities of multimodal models through a two-stage process: initial visual-only retrieval followed by a multimodal reranking stage. This methodology significantly improves the alignment between visual and textual information, leading to more accurate and contextually relevant answers. Experimentally, we have conducted thorough ablation studies to demonstrate the effectiveness of our proposed components. While comparing to existing state-of-the-art approaches on the Encyclopedic VQA and InfoSeek datasets, EchoSight demonstrates significant performance improvement, with an accuracy of 41.8% on E-VQA and 31.3% on InfoSeek. The success of EchoSight highlights the importance of efficient retrieval processes and the integration of multimodal information in enhancing the performance of large language models (LLMs) in knowledge-based VQA tasks.
Limitations
Although our proposed EchoSight demonstrates impressive performance on Knowledge-based VQA like Encyclopedic-VQA and InfoSeek, several limitations must be acknowledged. EchoSight’s performance is heavily dependent on the quality and comprehensiveness of the underlying knowledge base used for retrieval. Domain-specific knowledge not covered in these databases may lead to sub-optimal performance in specialized queries. In addition, the retrieval process, especially when involving multimodal reranking of candidates, introduces significant computational overheads, making it less suitable for real-time applications. These overheads can impact the efficiency and response time of the system. Future work focusing on improving the quality of knowledge bases and mitigating computational overheads remains to be explored.
Appendix A Dataset Details
In this section, we provide more details of in the Dataset we used. We summarize the statistics of in Table 9.
A.1 E-VQA
We focus only on Single-hop questions of E-VQA Mensink et al. (2023), namely Templated, Automatic, and Multi Answer questions in the table.
      Dataset       Question Type       Number of IQA pairs
      Train       Val       Test
      E-VQA       Templated       66,535       1,827       1,000
      Automatic       737,114       8,025       2,750
      Multi Answer       112,736       1,844       1,000
      Total       916,385       11,696       4,750
      InfoSeek       Total       902,509       -       71,335
Table 9: Dataset details used in our EchoSight’s traning and testing.
A.2 InfoSeek
And for Infoseek Chen et al. (2023), due to the missing entities in the knowledge-base we use, we remove the examples in the dataset. Specifically, 916,385 examples in training split out of 934,048 are kept (98.1%), and 71,335 examples of validation split out of 73,620 are kept (96.9%). Therefore, the results we obtain with our knowledge base are consistent with the dataset’s original setting while considering for the limitations of our knowledge base.
Appendix B Qualitative Results
Figure 3: Qualitative results showing the effective performance of our multimodal reranking method. While the reference image and question is displayed at left side, the visual-only search results are showcased at row 1 of each example, and the multimodal reranking results are showcased at row 2. The target image is squared with orange box.
B.1 reranking results
Qualitative results of our EchoSight’s multimodal reranking are as shown in Figure 3.
Figure 4: Qualitative VQA results from Encyclopedic VQA comparing to GPT-4V. The first row shows results in landmarks and the second row in natural species. Some failure cases are shown in the third row altogether with ground-truth.
B.2 VQA results
As shown in Figure 4, our EchoSight demonstrates significant improvements in multimodal understanding and generation tasks compared to the state-of-the-art GPT-4V Achiam et al. (2023).
Appendix C Prompt Template
C.1 E-VQA
The prompt we use for LLM when testing E-VQA Mensink et al. (2023) is shown as follow:
USER: Context: <CONTEXT>
Question: <QUESTION>
The answer is:
C.2 InfoSeek
Due to the strict metrics of exact match are used by InfoSeek Chen et al. (2023), we have to consider the format of the prompt so that the generated answer is comparable with the ground truth. Thereby, by using a one-shot example to keep the format correct, our prompt we use for InfoSeek is:
SYSTEM: You always answer the question the user asks. Do not answer anything else.
USER:Context: The sounthern side of the Alps is next to Lake Como.
Question: Which body of water is this mountain located in or next to?
Just answer the questions, no explanations needed.
Short answer is: Lake Como
Context: <CONTEXT>
Question: <QUESTION>
Just answer the questions, no explanations needed.
Short answer is: